# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ 3.5.4 Atomic quantum dots coupled to superfluid Bose-Einstein condensates
An alternative way, compared to the proposal of Ref. Felicetti et al. (2017a) in Sec. 3.5.3, to use cold-atom systems for the simulation of the quantum Rabi model was put forward in Ref. Felicetti et al. (2017b). In that proposal, a superfluid BEC in a shallow confining trap is made to interact with an atomic quantum dot Recati et al. (2005), i.e., a single atom (of the same species as the atoms in the BEC, but in a different hyperfine state) in a tight trap. The system is assumed to be in the collisional-blockade regime, meaning that the interaction $g_{dd}$ between atoms in the dot is much larger than interaction $g_{cc}$ between atoms in the BEC, such that the dot is never occupied by more than one atom, making it a two-level system. The BEC is assumed to be in the low-temperature superfluid regime, such that its quantum fluctuations can be described by a continuum of bosonic modes (phononic excitations). The coupling between the atomic quantum dot and the BEC is introduced through a Raman transition (effective Rabi frequency $\Omega$, detuning $\delta$) from external lasers.
@@
where $V$ is the volume of the BEC and $g_{cd}$ is the coupling between the atomic quantum dot and the atoms in the condensate. The parameter $\delta^{\prime}$ depends on $g_{cd}$, $g_{cc}$, $\Omega$, and $\delta$ Recati et al. (2005). Since the couplings $g_{cd}$, $g_{cc}$, and $g_{dd}$ can be tuned by external magnetic fields close to Feshbach resonances, it is possible to set $\delta^{\prime}=0$. By further having suitable boundary conditions for the BEC trap, the phononic modes can be spaced widely enough in frequency that only one of them effectively couples to the two-level system of the atomic quantum dot. We then recover the quantum Rabi Hamiltonian from Eq. (126).
In Ref. Felicetti et al. (2017b), it is shown that for standard parameters of a setup with a rubidium BEC and a potassium atomic quantum dot, one can achieve an effective cavity frequency of a few hundred Hz (clearly lower than in most other simulation proposals discussed in this review) and an effective coupling $g$ that ranges from much less than that to much more than that, i.e., spanning from much less than USC all the way to DSC. It is also possible to add more atomic quantum dots to simulate the Dicke model instead of the quantum Rabi model.
@@ 3.5.5 Ultrastrong coupling of two resonators via three-wave mixing
A close relative of the quantum Rabi and Dicke models is the Hopfield model Hopfield (1958), which describes the interaction between two harmonic modes (not a harmonic mode and a qubit; however, the second harmonic mode can be an effective model for a collection of atoms) without having made the RWA. Omitting the diamagnetic term, the Hopfield Hamiltonian reads
$H=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+G\left(a+a^{\dagger}\right)\left(b+b^{\dagger}\right),$ (167)
where $\omega_{a}$ and $\omega_{b}$ are the resonance frequencies of the modes $a$ and $b$, respectively, and $G$ is the strength of the coupling between them.
Figure 14: Setup of Ref. Fedortchenko et al. (2017) for simulating USC between two resonator modes, $a$ (orange) and $b$ (blue). The ring of four Josephson junctions shorted by inductors in the middle of the two resonators realizes three-wave mixing between $a$, $b$, and a mode $c$ that can be driven from the left in the setup. The box on the left is a hybrid coupler which serves to couple modes $a$ and $c$ to different transmission lines. Figure from Ref. Fedortchenko et al. (2017), S. Fedortchenko et al., Phys. Rev. A 95, 042313 (2017), with permission.
In 2017, Ref. Fedortchenko et al. (2017) put forward a proposal for simulating Eq. (167) in a setup with two superconducting resonators connected through a three-wave-mixing process generated by a Josephson mixer Abdo et al. (2013). By pumping the Josephson mixer with two tones, at frequencies $\omega_{B}=\omega_{a}+\omega_{b}+2\delta$ and $\omega_{R}=\omega_{a}-\omega_{b}$, respectively, where $\delta$ is a small detuning, the Hamiltonian for the setup (shown in Fig. 14) becomes
@@
The proposal of Ref. Fedortchenko et al. (2017) was implemented in an experiment in 2018 Marković et al. (2018). There, the mode frequencies were $\omega_{a}/2\pi=8.477\,\mathrm{GHz}$ and $\omega_{b}/2\pi=6.476\,\mathrm{GHz}$, much larger than the achievable interaction strength $\chi c_{B,R}$ from the three-wave mixing. However, the simulated effective frequency and coupling were of comparable size, with both a few tens of MHz. In this regime, Ref. Marković et al. (2018) demonstrated how the modes hybridize in the USC ground state, and measured both single- and two-mode squeezing of the emitted fields from the two resonators, as a result of the entanglement in the ground state.
@@ 3.5.6 Simulation of ultrastrong optomechanics
The USC regime in cavity optomechanics refers to the case where the single-photon optomechanical coupling $g_{0}$ is comparable to the mechanical frequency $\omega_{m}$, i.e., $g_{0}\sim\omega_{m}$. However, as mentioned in Sec. 2.1, the single-photon coupling $g_{0}$ is typically extremely weak, and reaching ultrastrong optomechanical interaction has remained challenging so far, despite considerable theoretical Rimberg et al. (2014); Nation et al. (2016); Shevchuk et al. (2017); Kounalakis et al. (2020); Manninen et al. (2022) and experimental Fogliano et al. (2021) effort aimed at this goal. Therefore, there have been several methods proposed to simulate ultrastrong optomechanical interaction with current technologies Johansson et al. (2014); Kim et al. (2015); Liao et al. (2020).
Among these possible simulation methods, it was suggested in Ref. Liao et al. (2020) that two bosonic modes, with annihilation operators $a$ and $b$, are coupled with a cross-Kerr interaction of strength $\chi$, i.e.,
@@
In an analogous way, it has been theoretically shown that a Fredkin-type interaction of three bosonic modes can also be used to simulate ultrastrong optomechanical interaction Yin et al. (2022), and that with the squeezing of the mode $b$, even a quadratic optomechanical coupling in the USC regime can be simulated Zhou et al. (2021).
@@ 4 Summary and perspectives
It should be stressed that the methods described here are not limited to photon-mediated interactions, but have also been applied to amplify interactions mediated by phonons or magnons. In fact, the first experimental demonstrations of quantum amplified boson-mediated interactions were reported for phonons Burd et al. (2021), and only later for photons Villiers et al. (2022). We believe that these described methods can also be generalized and applied to amplify the interactions mediated by exciton-polaritons, plasmons, magneto-plasmons, rotons, or other types of collective bosonic interactions.
@@
Despite this impressive progress, both theoretical and experimental, we hold the perspective that the exploration of the ultrastrong and deep-strong coupling regimes remains in its nascent stages, with a vast realm of novel effects and potential applications for quantum technologies awaiting their emergence. It should be stressed that such applications are not limited to quantum information processing, quantum metrology and sensing, advanced materials (like cavity-mediated electron-photon superconductors) and nanotechnology, but also include possible applications for quantum chemistry Levine et al. (2009), quantum biology Lambert et al. (2013), or for testing supersymmetric (SUSY) theories. We believe that, by applying the described or other methods for amplifying light-matter interactions, such fundamental and practical objectives can be accomplished with greater ease and speed.
@@ Acknowledgements
We thank Daniel H. Slichter and Marius Villiers for their valuable comments. W.Q. acknowledges support of the National Natural Science Foundation of China (NSFC) (Grants No. 0401260012 and No. 62131002). A.M. is supported by the Polish National Science Centre (NCN) under the Maestro Grant No. DEC-2019/34/A/ST2/00081. A.F.K. acknowledges support from the Japan Society for the Promotion of Science through a BRIDGE Fellowship, the Swedish Research Council (grant number 2019-03696), the Swedish Foundation for Strategic Research, the Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA via the project 101113946 OpenSuperQPlus100, and from the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). C.S.M. acknowledges support of a fellowship from “la Caixa” Foundation (ID 100010434) and from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 847648, with fellowship code LCF/BQ/PI20/11760026, and financial support from the MCINN project PID2021-126964OB-I00 (QENIGMA) and the Proyecto Sinérgico CAM 2020 Y2020/TCS-6545 (NanoQuCo-CM). F.N. is supported in part by: Nippon Telegraph and Telephone Corporation (NTT) Research, the Japan Science and Technology Agency (JST) [via the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant Number JPMJMS2061], the Asian Office of Aerospace Research and Development (AOARD) (via Grant No. FA2386-20-1-4069), and the Office of Naval Research (ONR).
@@ A Eliminating squeezing-induced noise
Following the standard procedure in Ref. Scully and Zubairy (1997) and then returning to the frame rotating at $H_{\rm rot}$, we obtain the master equation expressed, in terms of the $a_{s}$ mode, as
@@ B Effective Hamiltonians from perturbation theory
Our main goal is to obtain an effective Hamiltonian that describes the evolution of subsystem $A$ under the influence of subsystem $B$, without the need of explicitly describing subsystem $B$. This is only possible if the timescales of the intrinsic dynamics of the two subspaces are very different—this is why it is called adiabatic elimination—and still remain different even after considering the coupling between them. In energetic terms, this requires a “clustering” of energy levels, so that these energy levels can be split into some well-separated manifolds, as discussed, e.g., in Ref. Cohen‐Tannoudji et al. (1998) and illustrated in Fig. 1. In particular, by denoting the eigenvalues of $H_{A}$ and $H_{B}$ as $\{E_{A,i}\}$ and $\{E_{B,i}\}$, respectively, we require
@@ C Effective Hamiltonians in second-quantized form
In this section, we provide details of the derivation of the effective Hamiltonians, which describe the nonlinear characteristic of ultrastrong light-matter interactions simulated in JC-type systems with the addition of a single drive, as discussed in Sec. 3.3 of the main text. The three cases that we consider here are based on two examples of the cavity-QED systems depicted in Fig. 1: a single qubit coupled to two cavities, and two qubits coupled to a single cavity. The three nonlinear processes that we discuss are: (i) two photons simultaneously emitted by a single atom; (ii) frequency conversion of photons; and (iii) two atoms simultaneously excited by a single photon. These three cases are examples of nonlinear processes that have been discussed in the context of ultrastrongly coupled light-matter systems, e.g., in Refs. Garziano et al. (2016); Stassi et al. (2017); Kockum et al. (2017a).
@@ D.1 Example I: Two photons simultaneously emitted by a single atom
Figure 2: Intermediate transition processes that mediate the effective interactions (a) between the two states $\left|{+,n,m}\right\rangle$ and $\left|{-,n+1,m+1}\right\rangle$ in example I, (b) between the two states $\left|{+,n,m+1}\right\rangle$ and $\left|{-,n+1,m}\right\rangle$ in example II, and (c) between the two states $\left|{-,-,n+1}\right\rangle$ and $\left|{+,+,n}\right\rangle$ in example III. Blue solid arrows denote transitions that conserve the number of excitations, red dashed arrows denote transitions that change the number of excitations by one, and blue dashed arrows denote transitions that change the number of excitations by two. Transitions marked by blue arrows conserve parity; transitions denoted by red arrows do not. See also Refs. Garziano et al. (2016); Kockum et al. (2017a, b).
@@ D.3 Example III: Two atoms simultaneously excited by a single photon
Here we discuss the nonlinear effect originally described in the USC regime in Ref. Garziano et al. (2016) and recently experimentally realized in Ref. Tomonaga et al. (2023); that is, two atoms can be simultaneously excited by a single photon, and vice versa.

### v2

@@ 3.5.4 Atomic quantum dots coupled to superfluid Bose-Einstein condensates
An alternative way, compared to the proposal of Ref. [417] in Sec. 3.5.3, to use cold-atom systems for the simulation of the quantum Rabi model was put forward in Ref. [421]. In that proposal, a superfluid BEC in a shallow confining trap is made to interact with an atomic quantum dot [422], i.e., a single atom (of the same species as the atoms in the BEC, but in a different hyperfine state) in a tight trap. The system is assumed to be in the collisional-blockade regime, meaning that the interaction $g_{dd}$ between atoms in the dot is much larger than interaction $g_{cc}$ between atoms in the BEC, such that the dot is never occupied by more than one atom, making it a two-level system. The BEC is assumed to be in the low-temperature superfluid regime, such that its quantum fluctuations can be described by a continuum of bosonic modes (phononic excitations). The coupling between the atomic quantum dot and the BEC is introduced through a Raman transition (effective Rabi frequency $\Omega$, detuning $\delta$) from external lasers.
@@
where $V$ is the volume of the BEC and $g_{cd}$ is the coupling between the atomic quantum dot and the atoms in the condensate. The parameter $\delta^{\prime}$ depends on $g_{cd}$, $g_{cc}$, $\Omega$, and $\delta$ [422]. Since the couplings $g_{cd}$, $g_{cc}$, and $g_{dd}$ can be tuned by external magnetic fields close to Feshbach resonances, it is possible to set $\delta^{\prime}=0$. By further having suitable boundary conditions for the BEC trap, the phononic modes can be spaced widely enough in frequency that only one of them effectively couples to the two-level system of the atomic quantum dot. We then recover the quantum Rabi Hamiltonian from Eq. (126).
In Ref. [421], it is shown that for standard parameters of a setup with a rubidium BEC and a potassium atomic quantum dot, one can achieve an effective cavity frequency of a few hundred Hz (clearly lower than in most other simulation proposals discussed in this review) and an effective coupling $g$ that ranges from much less than that to much more than that, i.e., spanning from much less than USC all the way to DSC. It is also possible to add more atomic quantum dots to simulate the Dicke model instead of the quantum Rabi model.
@@ 3.5.5 Ultrastrong coupling of two resonators via three-wave mixing
A close relative of the quantum Rabi and Dicke models is the Hopfield model [32], which describes the interaction between two harmonic modes (not a harmonic mode and a qubit; however, the second harmonic mode can be an effective model for a collection of atoms) without having made the RWA. Omitting the diamagnetic term, the Hopfield Hamiltonian reads
$H_{\rm hpd}=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+G\left(a+a^{\dagger}\right)\left(b+b^{\dagger}\right),$ (171)
where $\omega_{a}$ and $\omega_{b}$ are the resonance frequencies of the modes $a$ and $b$, respectively, and $G$ is the strength of the coupling between them. Note that here, in order to satisfy the gauge principle, the Hopfield Hamiltonian in the dipole gauge is given by [406]
$H_{\rm hpd}^{\rm dg}=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b+iG_{\rm dg}(a^{\dagger}-a)(b+b^{\dagger})+G^{\prime}_{\rm dg}(b+b^{\dagger})^{2},$ (172)
and in the Coulomb gauge, by
$H_{\rm hpd}^{\rm dg}=\omega_{a}a^{\dagger}a+\omega_{b}b^{\dagger}b-iG_{\rm cg}(b^{\dagger}-b)(a+a^{\dagger})+G^{\prime}_{\rm cg}(a+a^{\dagger})^{2},$ (173)
where $G_{\rm dg}$ and $G_{\rm cg}$ are the coupling strengths in the dipole and Coulomb gauges, respectively. Moreover, $G^{\prime}_{\rm dg}$ and $G^{\prime}_{\rm cg}$ are the corresponding diamagnetic amplitudes, which are related to the strengths $G_{\rm dg}$ and $G_{\rm cg}$, respectively. The Hopfield Hamiltonian in Eq. (171) can be viewed as a simper variant in the dipole or Coulomb gauge, neglecting the diamagnetic term $\propto(b+b^{\dagger})^{2}$ or $(a+a^{\dagger})^{2}$. However, note that although these diamagnetic terms can be safely removed below the USC regime, they can become important in the USC regime [74, 406, 393].
Figure 14: Setup of Ref. [423] for simulating USC between two resonator modes, $a$ (orange) and $b$ (blue). The ring of four Josephson junctions shorted by inductors in the middle of the two resonators realizes three-wave mixing between $a$, $b$, and a mode $c$ that can be driven from the left in the setup. The box on the left is a hybrid coupler which serves to couple modes $a$ and $c$ to different transmission lines. Figure from Ref. [423], S. Fedortchenko et al., Phys. Rev. A 95, 042313 (2017), with permission.
In 2017, Ref. [423] put forward a proposal for simulating Eq. (171) in a setup with two superconducting resonators connected through a three-wave-mixing process generated by a Josephson mixer [424]. By pumping the Josephson mixer with two tones, at frequencies $\omega_{B}=\omega_{a}+\omega_{b}+2\delta$ and $\omega_{R}=\omega_{a}-\omega_{b}$, respectively, where $\delta$ is a small detuning, the Hamiltonian for the setup (shown in Fig. 14) becomes
@@
The proposal of Ref. [423] was implemented in an experiment in 2018 [425]. There, the mode frequencies were $\omega_{a}/2\pi=8.477\,\mathrm{GHz}$ and $\omega_{b}/2\pi=6.476\,\mathrm{GHz}$, much larger than the achievable interaction strength $\chi c_{B,R}$ from the three-wave mixing. However, the simulated effective frequency and coupling were of comparable size, with both a few tens of MHz. In this regime, Ref. [425] demonstrated how the modes hybridize in the USC ground state, and measured both single- and two-mode squeezing of the emitted fields from the two resonators, as a result of the entanglement in the ground state.
@@ 3.5.6 Simulation of ultrastrong optomechanics
The USC regime in cavity optomechanics refers to the case where the single-photon optomechanical coupling $g_{0}$ is comparable to the mechanical frequency $\omega_{m}$, i.e., $g_{0}\sim\omega_{m}$. However, as mentioned in Sec. 2.1, the single-photon coupling $g_{0}$ is typically extremely weak, and reaching ultrastrong optomechanical interaction has remained challenging so far, despite considerable theoretical [283, 286, 285, 287, 284] and experimental [426] effort aimed at this goal. Therefore, there have been several methods proposed to simulate ultrastrong optomechanical interaction with current technologies [427, 428, 429].
Among these possible simulation methods, it was suggested in Ref. [429] that two bosonic modes, with annihilation operators $a$ and $b$, are coupled with a cross-Kerr interaction of strength $\chi$, i.e.,
@@
In an analogous way, it has been theoretically shown that a Fredkin-type interaction of three bosonic modes can also be used to simulate ultrastrong optomechanical interaction [430], and that with the squeezing of the mode $b$, even a quadratic optomechanical coupling in the USC regime can be simulated [431].
@@ 4 Summary and perspectives
It should be stressed that the methods described here are not limited to photon-mediated interactions, but have also been applied to amplify interactions mediated by phonons or magnons. In fact, the first experimental demonstrations of quantum amplified boson-mediated interactions were reported for phonons [206], and only later for photons [209]. We believe that these described methods can also be generalized and applied to amplify the interactions mediated by exciton-polaritons, plasmons, magneto-plasmons, rotons, or other types of collective bosonic interactions.
@@
Despite this impressive progress, both theoretical and experimental, we hold the perspective that the exploration of the ultrastrong and deep-strong coupling regimes remains in its nascent stages, with a vast realm of novel effects and potential applications for quantum technologies awaiting their emergence. It should be stressed that such applications are not limited to quantum information processing, quantum metrology and sensing, advanced materials (like cavity-mediated electron-photon superconductors) and nanotechnology, but also include possible applications for quantum chemistry [25], quantum biology [26], or for testing supersymmetric (SUSY) theories. We believe that, by applying the described or other methods for amplifying light-matter interactions, such fundamental and practical objectives can be accomplished with greater ease and speed.
@@ Acknowledgements
We thank Daniel H. Slichter and Marius Villiers for their valuable comments. W.Q. acknowledges support of the National Natural Science Foundation of China (NSFC) (Grants No. 0401260012 and No. 62131002). A.M. is supported by the Polish National Science Centre (NCN) under the Maestro Grant No. DEC-2019/34/A/ST2/00081. A.F.K. acknowledges support from the Japan Society for the Promotion of Science through a BRIDGE Fellowship, the Swedish Research Council (grant number 2019-03696), the Swedish Foundation for Strategic Research, the Horizon Europe programme HORIZON-CL4-2022-QUANTUM-01-SGA via the project 101113946 OpenSuperQPlus100, and from the Knut and Alice Wallenberg Foundation through the Wallenberg Centre for Quantum Technology (WACQT). C.S.M. acknowledges support of a fellowship from “la Caixa” Foundation (ID 100010434) and from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 847648, with fellowship code LCF/BQ/PI20/11760026, and financial support from the MCINN project PID2021-126964OB-I00 (QENIGMA) and the Proyecto Sinérgico CAM 2020 Y2020/TCS-6545 (NanoQuCo-CM). F.N. is supported in part by: Nippon Telegraph and Telephone Corporation (NTT) Research, the Japan Science and Technology Agency (JST) [via the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant Number JPMJMS2061], the Asian Office of Aerospace Research and Development (AOARD) (via Grant No. FA2386-20-1-4069), and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074).
@@ A Eliminating squeezing-induced noise
Following the standard procedure in Ref. [1] and then returning to the frame rotating at $H_{\rm rot}$, we obtain the master equation expressed, in terms of the $a_{s}$ mode, as
@@ B Effective Hamiltonians from perturbation theory
Our main goal is to obtain an effective Hamiltonian that describes the evolution of subsystem $A$ under the influence of subsystem $B$, without the need of explicitly describing subsystem $B$. This is only possible if the timescales of the intrinsic dynamics of the two subspaces are very different—this is why it is called adiabatic elimination—and still remain different even after considering the coupling between them. In energetic terms, this requires a “clustering” of energy levels, so that these energy levels can be split into some well-separated manifolds, as discussed, e.g., in Ref. [373] and illustrated in Fig. 1. In particular, by denoting the eigenvalues of $H_{A}$ and $H_{B}$ as $\{E_{A,i}\}$ and $\{E_{B,i}\}$, respectively, we require
@@ C Effective Hamiltonians in second-quantized form
In this section, we provide details of the derivation of the effective Hamiltonians, which describe the nonlinear characteristic of ultrastrong light-matter interactions simulated in JC-type systems with the addition of a single drive, as discussed in Sec. 3.3 of the main text. The three cases that we consider here are based on two examples of the cavity-QED systems depicted in Fig. 1: a single qubit coupled to two cavities, and two qubits coupled to a single cavity. The three nonlinear processes that we discuss are: (i) two photons simultaneously emitted by a single atom; (ii) frequency conversion of photons; and (iii) two atoms simultaneously excited by a single photon. These three cases are examples of nonlinear processes that have been discussed in the context of ultrastrongly coupled light-matter systems, e.g., in Refs. [171, 85, 162].
@@ D.1 Example I: Two photons simultaneously emitted by a single atom
Figure 2: Intermediate transition processes that mediate the effective interactions (a) between the two states $\left|{+,n,m}\right\rangle$ and $\left|{-,n+1,m+1}\right\rangle$ in example I, (b) between the two states $\left|{+,n,m+1}\right\rangle$ and $\left|{-,n+1,m}\right\rangle$ in example II, and (c) between the two states $\left|{-,-,n+1}\right\rangle$ and $\left|{+,+,n}\right\rangle$ in example III. Blue solid arrows denote transitions that conserve the number of excitations, red dashed arrows denote transitions that change the number of excitations by one, and blue dashed arrows denote transitions that change the number of excitations by two. Transitions marked by blue arrows conserve parity; transitions denoted by red arrows do not. See also Refs. [171, 162, 163].
@@ D.3 Example III: Two atoms simultaneously excited by a single photon
Here we discuss the nonlinear effect originally described in the USC regime in Ref. [171] and recently experimentally realized in Ref. [174]; that is, two atoms can be simultaneously excited by a single photon, and vice versa.